OpenStudy (anonymous):

Solve the following system of equations: y=6x-11 -2x-3y=-7 Answer a. (2,4) b. (2, 1) c. (-3,1) d. (1, 2)

OpenStudy (anonymous):

substitute y=6x-11 in the 2º equation: -2x-3(6x-11)=-7 -2x-18x+33=-7 -20x=-40 x=2

OpenStudy (anonymous):

find y now

OpenStudy (anonymous):

can you show me?

OpenStudy (anonymous):

substitute value of x =2 in any equation and find y: y=6(2)-11=1

OpenStudy (anonymous):

@Tea-Is-God